# Social Optimization in Noncooperative Games under Central Regulation

### v1

@@ Abstract
This paper proposes a novel optimization problem building on noncooperative games under central regulation, which can be formulated as a bilevel structure. In the low-level, each player competes to minimize its own cost function that depends not only on the strategies of all players, but also on an intervention decision of the central regulator, while the central regulator located at the high-level attempts to achieve the social optimum, that is, to minimize the sum of cost functions of all players through an adjustable intervention decision. In this setting, under the intervention of the central regulator, the low-level players perform in a noncooperative game and aim to seek the Nash equilibrium, which indeed is related with the regulator’s decision. Meanwhile, the objective of the regulator is to choose a decision such that the social cost, i.e., the sum of cost functions of all players is minimum. This formulated bilevel social optimization problem is proven to be constrained, nonconvex and nonsmooth. To address this intricate problem, an inexact zeroth-order algorithm is developed by virtue of the smoothing techniques, allowing for the Nash equilibrium of the low-level game to be computed in an inexact manner. Levering the properties of smoothing techniques, it is rigorously shown that the devised algorithm achieves a sublinear convergence rate for computing a stationary point of a related optimization problem with a smoothed objective. Moreover, the sublinear convergence rate in the scenario where the exact equilibrium of the low-level game is available is also discussed. Finally, numerical simulations are conducted to demonstrate the efficiency of theoretical findings.
Index Terms:
@@
Social optimum, noncooperative games, nondifferentiable nonconvex optimization, smoothing techniques, gradient-free method.
I Introduction
@@
Game theory provides a suitable framework for modeling strategic interactions among rational players that aim to selfishly minimize their private, yet interdependent objective functions. This competitive setting has been commonly applied to various domains, including network congestion control [1], demand-side management in smart grids [2], and social networks [3]. One of most important issues in noncooperative games is to seek Nash equilibria (NEs) or generalized Nash equilibria (GNEs) which eventually dictate stable and satisfactory states wherein no player possesses an incentive for unilateral deviation.
@@
In the past decades, noncooperative games have been extensively explored under different scenarios. Generally speaking, the existing algorithms for seeking NEs of noncooperative games can be grouped into three categories. The simplest setup pertains to unconstrained games [4, 5], wherein each player possesses the liberty to select strategies from the entire real number or vector spaces. Moving towards a slightly more intricate case, games with local uncoupled decision set constraints are considered [6, 7], necessitating players to confine their decisions within predetermined nonempty sets, which are often presumed to be closed and convex. Furthermore, much attention is paid for games involving coupled equality or inequality constraints [8, 9].
@@
It should be noted that in noncooperative games, each player acts selfishly and only concerns its individual interest. However, such self-centered behaviors may lead to deterioration in performance of the social cost, that is, the sum of cost functions of all players. Specially, some existing studies have yielded valuable insights into the detrimental consequences of selfish behaviors on performance degradation in diverse fields, such as supply chains [10], resource allocation [11], and congestion games [12, 13] by virtue of the metrics of the price of anarchy and the price of stability [14, 15]. A prospective direction of research involves improving social cost while also accommodating the preferences of individual players in noncooperative games. Consequently, a central regulator has its purpose to regulate its decision to eusure that the social cost is minimum. For instance, in the context of economic investment problems within network games, the central blank, who wishes to ensure market stability and avoid systemic risk, maximizes the overall social return for all participants by meticulously regulating individual standalone marginal return [16]. Similarly, in power systems, the total social resource consumption is minimized through the strategic adjustment of electricity prices [17].
@@
To date, there has been some literature on adaptive incentive mechanisms for steering NEs towards a desirable point [18, 19, 20, 21, 22]. However, these works were explored from the control theoretic perspective, and paid attention to sufficient conditions or direct computation of the best targeting intervention for problems exhibiting specific structures. Recently, combining with the gradient dynamics, the authors in [16] studied continuous-time intervention protocols based on limited knowledge, which were rigorously shown to successfully drive actions of players in quadratic network games to the social optimum. For nonconvex cooperative aggregative games, a distributed stochastic annealing algorithm was developed in [23] without external intervention and was proved to weakly converge. Note that both of [16] and [23] have not yet derived any result regarding the convergence rate. Furthermore, most of the aforementioned works did not consider games with set constraints and intervention constraints. In [24], a bilevel incentive design algorithm is devised for probability simplex constrained games and the sublinear rate of convergence is provided.
@@
It is emphasized that the exploration of social optima in noncooperative games is devoted to the global objective of the social optimization problem, thereby facilitating the central regulator in further improving the resource allocation efficiency. From a theoretical standpoint, when players’ objective functions hinge on their own decisions and those of other players, as well as a decision determined by the central regulator. Then, the corresponding situation extends beyond the scope of traditional noncooperative games, shifting the focus towards achieving a global optimal objective. This problem becomes increasingly complex and challenging due to its inherent bilevel structure, particularly, the mutual coupling of decisions between players and the central regulator. Therefore, the problem of achieving the social optimum in general noncooperative games remains to be an open and challenging issue.
@@
Motivated by the above analysis, to overcome the obstacle that the behaviors of selfish players proposing to minimize their own costs may lead to that the social cost is not optimal, this paper investigates the social optimization problem, which inherently features a bilevel structure. In the low-level, players participate in a noncooperative game, where each individual cost function relies not only on all players’ strategies but also on an intervention decision of a central regulator. Under the intervention of the regulator, all players endeavor to seek the NE corresponding to the regulator’s decision. Then, in the high-level, the central regulator tries to find an optimal decision so as to achieve the social optimum. The significant challenges arise from the absence of a closed-form expression for the low-level game, as well as the inherent nondifferentiability, nonconvexity, and constraints of the high-level function. An inexact gradient-free method by utilizing smoothing techniques is devised to solve this problem and elaborately analyzed to exhibit sublinear convergence rate. The main contributions are summarized as follows.
@@
    A new social optimization framework based on noncooperative games is formulated for the first time which is inherently equipped with a bilevel structure, where the strategies of low-level players and the decisions of the high-level regulator are intricately and mutually coupled. Although the related problems regarding social optimization were investigated in [23, 16, 24], the problem settings are rather different. Specifically, there was no intervention involved in [23]. In fact, players are reluctant to sacrifice their personal wellbeing to move toward the social optimum without external intervention. Moreover, to design interventions for special quadratic network games [16], the regulator is presumed to be aware of the social optimum in advance. However, this is often difficult or even impossible in reality. Therefore, our concerned problem is more general and practical.
@@
    The formulated problem is indeed constrained, nonsmooth, and nonconvex, leading to difficulties in problem solving. By employing the randomized smoothness and Moreau smoothness techniques, the studied problem is transformed into a corresponding smoothed unconstrained optimization problem. Then a zeroth-order algorithm is proposed, which enables the low-level game to be solved in an inexact manner to contend with the unavailable exact equilibrium of the low-level game. The devised algorithm utilizes random gradient-free oracles instead of relying on direct gradient information, and does not require the exact equilibrium of the low-level game. As a result, the algorithm becomes more practical and flexible in scenarios where obtaining gradients of the objective functions is infeasible or the exact NE is prohibitively expensive to evaluate.
@@
    Through rigorous analysis, in the case where the central regulator cannot obtain the NE of the low-level game, the sublinear convergence rate is established for computing a stationary point of the corresponding smoothed unconstrained optimization problem within the inexact regime. Furthermore, the sublinear rate is also derived in the scenario where the exact NE of the low-level game is available. However, the convergence rate of algorithms for social optimization in [16, 23] was not established.
The rest of this paper is outlined as follows. Section II formulates the problem and introduces the preliminaries. A zeroth-order algorithm is provided in Section III, followed by the convergence results. In Section IV, numerical simulations are given to illustrate the theoretical findings. Section V makes a brief conclusion.
@@
II PROBLEM FORMULATION AND PRELIMINARIES
II-A Problem Formulation
This paper introduces a novel social optimization framework grounded in noncooperative games consisting of $N$ players located at a low-level and a central regulator in the high-level. In the low-level, players perform in a noncooperative game and aim to selfishly minimize their own cost functions, dependent on not only the strategies of all players but also an adjustable intervention decisions of the central regulator. When the central regulator makes a decision, all players response at once simultaneously, aiming to seek the NE, which indeed is related to the regulator’s decision. Then the central regulator in the high-level seeks to determine an optimal intervention decision to reach the social optimum. Specifically, the schematic sketch of the problem setup is presented in Fig. 1, and the further expressions are detailed below.
Figure 1: Sketch of the problem setup. Each of the $N$ players in the low-level aims to minimize the personal cost function $f_{i}$ given intervention decision $\theta$ of the central regulator through a communication network. The regulator in the high-level utilizes required data received from all players to choose an optimal intervention decision to achieve social optimum. Here, $x_{-i}:=\operatorname*{col}\{x_{i}\}_{i\in[N]\backslash\{i\}}$ and $x(\theta):=\operatorname*{col}\{x_{i}(\theta)\}_{i\in[N]}$.
Low-level game. Each player $i\in[N]$ takes strategy $x_{i}$ from its strategy set $X_{i}\subseteq\mathbb{R}^{d_{i}}$. Let $x=\operatorname*{col}\{x_{i}\}_{i\in[N]}\in X:=X_{1}\times\cdots\times X_{N}\subseteq\mathbb{R}^{d}$ with $d=\sum_{i=1}^{N}d_{i}$ and $x_{-i}=\operatorname*{col}\{x_{i}\}_{i\in[N]\backslash\{i\}}\in{X_{-i}}:=X_{1}\times\cdots\times X_{i-1}\times X_{i+1}\times\cdots\times X_{N}\subseteq\mathbb{R}^{d-d_{i}}$ denote, respectively, the joint decision of all players and the joint strategy of all other players except $i$. The cost function of player $i$ is $f_{i}(x_{i},x_{-i},\theta)$ which depends on $x_{i}$, $x_{-i}$ and the central regulator’s intervention decision $\theta$ selected from a feasible set $\Theta\subseteq\mathbb{R}^{n}$. Herein, each player is not willing to share its own strategy and objective function with other players for the reason of privacy, and instead only has access to the information of partial players through local information exchange. Given a current decision $\theta$, players select their strategies with the aim of minimizing individual objective functions $f_{i}(x_{i},x_{-i},\theta)$, $i\in[N]$ to seek the NE that is in their best interest. This leads to a collection of interdependent optimization problems as follows:
@@ II-A Problem Formulation
In this study, we limit our focus to the case where the low-level game (1) admits a unique NE for any particular parametrization $\theta\in\Theta$. Upon completion of the game process, an NE $x(\theta)=(x(\theta),x_{-i}(\theta))$ is obtained, which corresponds to a set of cost functions of players: $f_{i}(x_{i}(\theta),x_{-i}(\theta),\theta),i\in[N]$.
@@
For any central regulator’s decision $\theta\in\mathbb{R}^{n}$, a strategy profile $x(\theta)=(x_{i}(\theta),x_{-i}(\theta))\in X$ is an NE for game (1) if for every player $i\in[N]$, it holds that
@@
High-level optimization. After receiving all players’ feedback $x_{i}(\theta)$, $i\in[N]$, the central regulator endeavors to determine the optimal decision $\theta^{*}$ that minimizes the sum of cost functions of all players, thereby achieving the social optimum. More precisely, the regulator aims to solve the following optimization problem:
$\min_{\theta\in\Theta}f(x(\theta),\theta):=\sum_{i=1}^{N}f_{i}(x(\theta),\theta).$ (3)
Note that the information on the players is available to the central regulator, making it preferable to leverage required data for effective coordination and decision-making.
In conclusion, the studied bilevel social optimization problem can be formulated as
@@
[…]
Consequently, the main objective of this paper is to seek the optimal decision $\theta^{*}$ to ultimately achieve the social optimum.
@@
Note that the key distinction between the social optimization (II-A) and Stackelberg games studied in [25, 26, 27] lies in the role of the high-level leader. Specifically, in the realm of Stackelberg games, the leader pursues the minimization of its own objective while adhering to the equilibrium constraints established by the low-level game among the followers. However, the central regulator in social optimization is tasked with optimizing the social cost. Moreover, in noncooperative games, only players compete with each other to minimize their own cost functions and no other regulators are involved. Therefore, the studied problem is essentially different from these frameworks since it is driven by the overall benefit to society, with a central regulator coordinating the whole network to reach this goal.
Some standard assumptions are listed next.
@@
    The set $X_{i}$ for any $i\in[N]$ is nonempty, compact, and convex.
@@
From Assumption 1-i) and iii), it can be further obtained that for any $i\in[N]$, $\theta\in\mathbb{R}^{n}$, and $x,x^{\prime}\in X$, there exist $B_{X},L_{x}>0$ such that
$\|x\|\leq B_{X},\ \|f_{i}(x,\theta)-f_{i}(x^{\prime},\theta)\|\leq L_{x}\|x-x^{\prime}\|.$ (6)
@@
    $(G(x,\theta)-G(x^{\prime},\theta))^{\textit{T}}(x-x^{\prime})\geq\mu\|x-x^{\prime}\|^{2},$ (8)
@@
Under Assumption 1-i) and iii), it follows from Proposition 1.4.2 in [28] that for any given $\theta\in\mathbb{R}^{n}$, a solution $x(\theta)\in X$ is an NE of game (1) if and only if
@@
The existence of the NE is guaranteed from Corollary 2.2.5 in [28]. Additionally, Assumption 2-i) ensures the uniqueness of the NE by Theorem 2.3.3 in [28], which is also made in [7, 29, 4, 6, 9].
Lemma 1.
Under Assumptions 1 and 2, $f_{i}(x(\theta),\theta)$ is Lipschitz continuous, i.e., there exists a constant $L_{F}>0$ satisfying
Proof.
See Appendix A. ∎
The communication pattern among players is characterized by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{A})$, where $\mathcal{V}=[N]$ is the player set, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the directed edge set, and $\mathcal{A}=[a_{ij}]\in\mathbb{R}^{N\times N}$ is the adjacency matrix. $a_{ij}>0$ if $(j,i)\in\mathcal{E}$ or $i=j$, and $a_{ij}=0$, otherwise. Here, a directed edge $(j,i)\in\mathcal{E}$ means player $j$ could send information to player $i$. Let $\mathcal{N}_{i}:=\{j\in[N]:(j,i)\in\mathcal{E}\}$ be the set of neighbors of player $j$. A directed path from $i_{1}$ to $i_{s}$ is defined as a sequence of edges $(i_{p},i_{p+1})\in\mathcal{E},p\in[s-1]$ for different nodes. A directed graph is said to be strongly connected if there is a directed path from every node to any other node. It is assumed that $a_{ii}>0$ for all $i\in[N]$ in this paper. In the following, a standard condition on players’ interaction is imposed.

### v2

@@ Abstract
Motivated by the increasing attention to overall social benefits in networked multi-agent systems, this paper investigates an optimization problem building on noncooperative games under high-level regulation, which can be formulated in a bilevel structure. Specifically, the low level consists of a noncooperative game, where each player competes to minimize its own cost function that depends not only on the strategies of all players, but also on an intervention decision of a regulator located at the high level. Under the intervention of the high-level regulator, the low-level players aim to seek a Nash equilibrium (NE), which indeed is related to the regulator’s decision. Meanwhile, the regulator in the high level attempts to achieve the social optimum, that is, to minimize the sum of all players’ costs obtained at the NE. This bilevel social optimization problem is proven to be nonconvex and nonsmooth, leading to challenges for solving it effectively, as the exact gradient of cost sum functions may not be available. To address this intricate problem, an inexact zeroth-order algorithm is developed by virtue of the smoothing techniques, allowing for approximating the NE of the low-level game and thus estimating the required gradients. It is rigorously shown that the devised algorithm achieves a sublinear convergence rate for computing an approximate stationary point of the studied problem. Moreover, the sublinear convergence rate in the scenario where the exact equilibrium of the low-level game is available is established. Finally, numerical simulations are conducted to demonstrate the efficiency of theoretical findings.
keywords:
@@
Social optimum, noncooperative games, nonsmooth nonconvex optimization, smoothing techniques, gradient-free method.
journal: Automatica
1 Introduction
@@
Game theory provides a suitable framework for modeling strategic interactions among rational players that aim to selfishly minimize their private, yet interdependent objective functions. This competitive setting has been commonly applied to various domains, including power allocation [sun25Learning], network congestion control [barrera2014dynamic], imitation and reinforcement learning [Hierarchical23], and social networks [mani21quan]. One of most important issues in noncooperative games is to seek Nash equilibria (NEs) or generalized Nash equilibria (GNEs) which eventually dictate stable and satisfactory states wherein no player possesses an incentive for unilateral deviation.
@@
In the past decades, noncooperative games have been extensively explored under different scenarios. Generally speaking, the existing algorithms for seeking NEs of noncooperative games can be grouped into three categories. The simplest setup pertains to unconstrained games [Nguyen23], wherein each player possesses the liberty to select strategies from the entire real number or vector spaces. Moving towards a slightly more intricate case, games with local uncoupled decision set constraints are considered [FengRR24, Nguyen25, MENG2023110919, Bianchi21], necessitating players to confine their decisions within predetermined nonempty sets, which are often presumed to be closed and convex. Furthermore, much attention is paid for games involving coupled equality or inequality constraints [meng2022linear, Gadjov21, meng24online, Carnevale2022Tracking-based]. An alternative perspective categorizes methods for seeking NE or GNE into two groups, according to the way how information is exchanged. Specifically, one class is the full-decision information scenario, where each player has access to the decisions taken by all others [belgioioso2018, lei22, Persis20]. Unfortunately, in such a scenario, a central coordinator is typically required to broadcast the data to the network, which is impractical in numerous real-world applications [belgioioso2018, ghaderi2014opinion]. The other case that has gained attention recently is the partial-decision information scenario, where each player is capable of estimating decisions of others based only on local information exchanges with neighbors through a communication network [Nguyen23, MENG2023110919, FengRR24, meng2022linear, Gadjov21, Bianchi21, Nguyen25, meng24online, Carnevale2022Tracking-based], resulting in fully distributed NE seeking algorithms.
@@
It should be noted that in noncooperative games, each player acts selfishly and only concerns its individual interest. However, such self-centered behaviors may lead to deterioration in performance of the social cost, that is, the sum of cost functions of all players. Specially, some existing studies have yielded valuable insights into the detrimental consequences of selfish behaviors on performance degradation in diverse fields, such as supply chains [ye16chain], resource allocation [johari05], federated edge learning [pan25bilateral], and congestion games [Roughgarden02, apt14] by virtue of the metrics of the price of anarchy and the price of stability [Cominetti24, Anshelevich08, Chandan24metho]. A prospective direction of research involves improving social cost while also accommodating the preferences of individual players in noncooperative games. Consequently, a high-level regulator has its purpose to regulate its decision to eusure that the social cost is minimum. For instance, in the context of economic investment problems within network games, the central bank, who wishes to ensure market stability and avoid systemic risk, maximizes the overall social return for all participants by meticulously regulating individual standalone marginal return [Shakarami23]. Similarly, in power systems, the total social resource consumption is minimized through the strategic adjustment of electricity prices [Ma14, sun25Learning].
@@
To date, there has been some literature on adaptive incentive mechanisms for steering NEs towards a desirable point [Galeotti2020, alpcan09, Maheshwari2021DynamicTF, ratliff21, Maheshwari22]. However, these works were explored from the systems and control theoretic perspective, and paid attention to sufficient conditions or direct computation of the best targeting intervention for problems exhibiting specific structures. Moreover, incentive designs for Stackelberg games were investigated in [bacsar2024incentive, Shen2007, sanjari2025incentive], where leaders attempt to shape the behavior of players by providing incentives. Notably, the leader in such settings are self-interested, as it pursues only its own objective and the optimal performance is ultimately attained from the leader’ viewpoint, simultaneously leading to an NE among players. This motivation is often driven by personal interests rather than by the concern of social benifits. Recently, combining with the gradient dynamics, the authors in [Shakarami23] studied continuous-time intervention protocols based on limited knowledge, which were rigorously shown to successfully drive actions of players in quadratic network games to the social optimum. For social optimization in nonconvex cooperative aggregative games, a distributed stochastic annealing algorithm was developed in [wang22] without external intervention and was proved to weakly converge. Note that both of [Shakarami23] and [wang22] have not yet derived any result regarding the convergence rate. Furthermore, most of the aforementioned works did not consider games with set constraints and intervention constraints. In [liu22], a bilevel incentive design algorithm is devised for probability simplex constrained games and the sublinear convergence rate is provided.
@@
It is emphasized that the exploration of social optima in noncooperative games is toward the global objective of a social optimization problem, thereby facilitating the high-level regulator in further improving resource allocation efficiency. From a theoretical viewpoint, when each player’s cost function depends on its own strategy and those of other players, as well as a decision determined by the high-level regulator, the resulting situation extends beyond the scope of traditional noncooperative games, shifting the focus towards attaining a global optimal objective. The problem becomes increasingly complex and challenging due to its inherent bilevel structure, particularly, the mutual coupling of decision-making process between players and the high-level regulator. Therefore, achieving the social optimum in general noncooperative games remains an open and challenging issue.
@@
Motivated by the above analysis, to overcome the obstacle that the behaviors of selfish players proposing to minimize their own costs may lead to the social cost not optimal, this paper investigates the social optimization problem, which inherently features a bilevel structure. In the low level, players participate in a noncooperative game, where each individual cost function relies not only on all players’ strategies but also on an intervention decision of a high-level regulator. Under the intervention of the regulator, all players endeavor to seek the NE corresponding to the regulator’s decision. Then, in the high level, the regulator tries to find an optimal decision so as to achieve the social optimum. The significant challenges arise from the lack of a closed-form solution for the low-level game, along with the inherent nondifferentiability and nonconvexity of the high-level function. An inexact gradient-free method by utilizing smoothing techniques is devised to solve this problem and elaborately analyzed to exhibit sublinear convergence rate. The main contributions are summarized as follows.
@@
    A new social optimization framework based on noncooperative games is formulated which is inherently equipped with a bilevel structure, where the strategies of low-level players and the decisions of the high-level regulator are intricately and mutually coupled. Although the related problems regarding social optimization were investigated in [wang22, Shakarami23, liu22], the problem settings are rather different. Specifically, there was no intervention involved in [wang22]. In fact, players are reluctant to sacrifice their personal wellbeing to move toward the social optimum without external intervention. Moreover, to design interventions for special quadratic network games [Shakarami23], the regulator is presumed to be aware of the social optimum in advance. However, this is often difficult or even impossible in reality. Therefore, our concerned problem is more general and practical.
@@
    The formulated problem is indeed constrained, nonsmooth, and nonconvex, leading to difficulties in problem solving. This is also the exact gradient of cost sum functions may not be available. By employing the randomized smoothness and Moreau smoothness techniques, the studied problem is transformed into a corresponding smoothed unconstrained optimization problem in form. Then a zeroth-order algorithm is proposed, which enables the low-level game to be solved in an inexact manner to contend with the unavailable exact equilibrium of the low-level game. The devised algorithm utilizes random gradient-free oracles instead of relying on direct gradient information, and does not require the exact equilibrium of the low-level game. As a result, the algorithm becomes more practical and flexible in scenarios where obtaining gradients of the objective functions is infeasible or the exact NE is prohibitively expensive to evaluate.
@@
    Through rigorous analysis, in the case where the high-level regulator cannot obtain the NE of the low-level game, the sublinear convergence rate is established for computing an approximate stationary point of the concerned problem within the inexact regime. Furthermore, the sublinear rate is also derived in the scenario where the exact NE of the low-level game is available. However, the convergence rate of algorithms for social optimization in [Shakarami23, wang22] was not established.
The rest of this paper is outlined as follows. Section 2 formulates the problem and introduces the preliminaries. A zeroth-order algorithm is provided in Section 3, followed by the convergence results. In Section 4, numerical simulations are given to illustrate the theoretical findings. Section 5 makes a brief conclusion.
Notations: $\mathbb{R}^{n}$ and $\mathbb{R}^{m\times n}$ denote the sets of $n$-dimensional real column vectors and $m\times n$ real matrices, respectively, endowed with Euclidean norm $\|\cdot\|$. Denote by $M^{\top}$ the transpose of a matrix or vector $M$. $I_{N}\in\mathbb{R}^{N\times N}$ represents the $N\times N$ identity matrix. $\bm{1}_{N}:=(1\ 1\ldots 1)^{\top}\in\mathbb{R}^{N}$. $\operatorname*{col}\{x_{i}\}_{i=1}^{N}:=(x_{1}^{\top},\ldots,x_{N}^{\top})^{\top}$. For an integer $m>0$, $m]:=\{1,\ldots,m\}$. $\otimes$ denotes the Kronecker product. For two matrices $P,Q$, $P\succ Q$ represents PQP-Q is positive definite. $\lceil a\rceil$ is the smallest integer greater than or equal to scalar $a$. Denote by $\mathrm{E}[\cdot]$ the expectation of a random variable. Given a closed convex set $C\subseteq\mathbb{R}^{n}$, $\Pi_{C}(x):=\arg\min_{y\in C}\|x-y\|$ stands for the projection of $x\in\mathbb{R}^{n}$ onto $C$, $\operatorname*{dist}(x,C):=\min_{y\in C}\|x-y\|$ denotes the distance of $x$ to $C$, and $\mathrm{N}_{C}(x):=\{s\in\mathbb{R}^{n}:s^{\top}(y-x)\leq 0,\forall y\in C\}$ represents the normal cone of $x\in C$ at $C$. For real-valued sequences $\{a_{k}\}$ and $\{b_{k}\}$, $a_{k}=\mathcal{O}(b_{k})$ if there exists $C>0$ satisfying $a_{k}\leq Cb_{k}$. Let $f:\mathbb{R}^{n}\to(-\infty,\infty]$ be a proper function, the Clarke generalized gradient of $f$ at $x$ is denoted as $\partial f(x):=\{w\in\mathbb{R}^{n}:f^{\circ}(x,d)\geq w^{\top}d,\forall d\in\mathbb{R}^{n}\}$, where $f^{\circ}(x,d):=\limsup_{z\to x,t\downarrow 0}(f(z+td)-f(z))/t$ is the generalized directional derivative of $f$ at $x$ in a given direction $d$. Let $\mathbb{S}^{n}:=\{u\in\mathbb{R}^{n}:\|u\|=1$} and $\mathbb{B}^{n}:=\{\nu\in\mathbb{R}^{n}:\|\nu\|\leq 1\}$ be $n$-dimensional unit ball and its surface, respectively. $\sigma_{N-1}(A)$ denotes the second largest singular value of matrix $A\in\mathbb{R}^{N\times N}$. For a function $f=(f_{1},f_{2},\ldots,f_{d})$ with $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ for $i=1,2,\ldots,n$, denote $\nabla f(x):=[\nabla f_{1}(x),\ldots,\nabla f_{d}(x)]$.
2 Problem formulation and preliminaries
2.1 Problem formulation
This paper introduces a novel social optimization framework grounded in noncooperative games consisting of $N$ players located at low level and a regulator in the high level. In the low level, each player performs in a noncooperative game and aims to selfishly minimize its own cost function, dependent on not only the strategies of all players but also an adjustable intervention decision of the high-level regulator. After the high-level regulator makes a decision first, all players response at once simultaneously, aiming to seek the NE, which indeed is related to the regulator’s decision. Then the regulator in the high level seeks to determine an optimal intervention decision to reach the social optimum. Specifically, the schematic sketch of the problem setup is presented in Fig. 1, and the further expressions are detailed below.
@@ 2.1 Problem formulation
[…]
@@
In this study, we limit our focus to the case where the low-level game (1) admits a unique NE for any particular parametrization $\theta\in\Theta$. Upon completion of the game process, an NE $x(\theta)=(x(\theta),x_{-i}(\theta))$ is obtained and defined in the following, which corresponds to a set of costs of players: $f_{i}(x_{i}(\theta),x_{-i}(\theta),\theta),i\in[N]$.
@@
For any $\theta\in\mathbb{R}^{n}$, a strategy profile $x(\theta)=(x_{i}(\theta),x_{-i}(\theta))\in X$ is an NE for game (1) if for every player $i\in[N]$, it holds that
@@
High-level optimization. After receiving all players’ feedback $x_{i}(\theta)$, $i\in[N]$, where $x(\theta)$ denotes the NE of the low-level game (1), the high-level regulator endeavors to determine the optimal decision $\theta^{*}$ that minimizes the sum of cost functions of all players, thereby achieving the social optimum.
Although the information on the players is available to the high-level regulator, the decisions of the regulator dynamically depend on the NE of the low-level game, resulting in high complexity and strong coupling. The goal of this paper is to address the following bilevel social optimization problem
@@
$\text{s.t.}~f_{i}(x_{i}(\theta),x_{-i}(\theta),\theta)\leq f_{i}(x_{i},x_{-i}(\theta),\theta),$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\ \forall i\in[N],x_{i}\in X_{i}.$ (3)
The studied bilevel social optimization probem is inherently nonconvex and nonsmooth, even when each $f_{i}$ is continuously differentiable and (strong) convex with respect to all its arguments, demonstrated by the following example.
Example 1.
Consider a noncooperative game with $2$ players. The cost function of player $i$ is $f_{i}(x_{i},x_{-i},\theta)=x_{i}^{2}-2x_{-i}-2x_{i}\theta$ and the strategy set is $X_{i}=[2/3,1]$, $\forall i\in[2]$. The feasible set of the high-level regulator is $\Theta=[0,1]$. Based on the optimal condition (11), it can be derived that
Through simple calculation, one has
$f(x(\theta),\theta)$ $=\sum_{i=1}^{2}f_{i}(x(\theta),\theta)$
$=\left\{\begin{aligned} &-\frac{4}{3}\theta-\frac{8}{9},&\text{if}\ \theta\in[0,2/3],\\ &-2\theta^{2}-4\theta,&\text{if}\ \theta\in[2/3,1].\end{aligned}\right$.
@@
Note that the key distinction between the social optimization (2.1) and Stackelberg games studied in [Maljkovic2024OnDC, linon24, bacsar2024incentive, Shen2007, sanjari2025incentive] lies in the role of the high-level leader. Specifically, in the realm of Stackelberg games, the leader pursues the minimization of its own objective while adhering to the equilibrium constraints established by the low-level game among the followers. However, the high-level regulator in social optimization is tasked with optimizing the social cost. Moreover, in noncooperative games, only players compete with each other to minimize their own cost functions and no other regulators are involved. Actually, both leaders and players in Stackelberg games and noncooperative games are selfish, as they primarily focus on self-interested behaviors. Therefore, the studied problem is essentially different from these frameworks since it is driven by the overall benefit to society, with a high-level regulator coordinating the whole network to reach this goal.
Some standard assumptions are listed next, which are commonly made in existing literature [meng2022linear, Nguyen23, FengRR24, qu24online, Bianchi21, Carnevale2022Tracking-based].
@@
    The set $X_{i}$ for any $i\in[N]$ is nonempty, compact and convex.
@@
From Assumption 1-i) and iii), it can be further obtained that for any $i\in[N]$, $\theta\in\mathbb{R}^{n}$, and $x,x^{\prime}\in X$, there exist positive constants $B_{X}$ and $L_{x}$ such that
$\|x\|\leq B_{X},$ (5)
$\|f_{i}(x,\theta)-f_{i}(x^{\prime},\theta)\|\leq L_{x}\|x-x^{\prime}\|.$ (6)
@@
    $(G(x,\theta)-G(x^{\prime},\theta))^{\top}(x-x^{\prime})\geq\mu\|x-x^{\prime}\|^{2},$ (8)
@@
Under Assumption 1-i) and iii), it follows from Proposition 1.4.2 in [facchinei2003finite] that for any given $\theta\in\mathbb{R}^{n}$, a solution $x(\theta)\in X$ is an NE of game (1) if and only if
@@
The existence of the NE is guaranteed from Corollary 2.2.5 in [facchinei2003finite]. Additionally, Assumption 2-i) ensures the uniqueness of the NE by Theorem 2.3.3 in [facchinei2003finite].
Remark 2.
Assumption 2 holds in numerous real-world scenarios. For instance, $f_{i}(x_{i},x_{-i},\theta)$ takes the form of $g_{i}(x_{i},x_{-i})+x_{i}^{\top}Q_{i}\theta$. Assume that the pseudo-gradient of $g_{i}(x_{i},x_{-i})$ is $\mu$-strongly monotone and $l$-Lipschitz, which are rather common in existing literature including economic investment and smart grids fields [Shakarami23, Maheshwari22, Maljkovic23].
The communication graph among low-level players is characterized by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{A})$, where $\mathcal{V}=[N]$ is the player set, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the directed edge set, and $\mathcal{A}=[a_{ij}]\in\mathbb{R}^{N\times N}$ is the adjacency matrix. $a_{ij}>0$ if $(j,i)\in\mathcal{E}$ or $i=j$, and $a_{ij}=0$, otherwise. Here, a directed edge $(j,i)\in\mathcal{E}$ means player $j$ could send information to player $i$. Let $\mathcal{N}_{i}:=\{j\in[N]:(j,i)\in\mathcal{E}\}$ be the set of neighbors of player $j$. A directed path from $i_{1}$ to $i_{s}$ is defined as a sequence of edges $(i_{p},i_{p+1})\in\mathcal{E},p\in[s-1]$ for different nodes. A directed graph is said to be strongly connected if there is a directed path from every node to any other node. It is assumed that $a_{ii}>0$ for all $i\in[N]$ in this paper. In the following, a standard condition on players’ interaction is imposed.